# From Landau two-fluid model to de Sitter Universe

### v10

@@ Abstract
The condensed matter analogs are useful for consideration of the phenomena related to the quantum vacuum. This is because in condensed matter we know physics both in the infrared and in the ultraviolet limits, while in particle physics and gravity the physics at trans-Planckian scale is unknown. One of the corner stones of the connections between the non-relativistic condensed matter and the modern relativistic theories is the two-fluid hydrodynamics of superfluid helium, which was developed by Landau and Khalatnikov (Isaak Markovich Khalatnikov later founded and headed the Landau Institute). The dynamics and thermodynamics of the de Sitter state of the expansion of the Universe bear some features of the multi-fluid system. There are actually three components: the quantum vacuum, the gravitational component and relativistic matter. The expanding de Sitter vacuum serves as the thermal bath with local temperature, which is twice the Gibbons-Hawking temperature related to the cosmological horizon. This local temperature leads to the heating of matter component and the gravitational component. The latter behaves as Zel’dovich stiff matter and represents the dark matter. In equilibrium and in the absence of the conventional matter the positive partial pressure of dark matter compensates the negative partial pressure of quantum vacuum. That is why in the full equilibrium the total pressure is zero. This is rather similar to the correspondingly superfluid and normal components of superfluid liquid, which together produce the zero pressure of the liquid in the absence of environment. We suggest the phenomenological theory, which describes the dynamics of the dark energy and the dark matter. If one assumes that in dynamics, the gravitational dark matter behaves as the real Zel’dovich stiff matter, one obtains that both components experience the power law decay due to the energy exchange between these components. Then it follows that their values at present time have the correct order of magnitude. We also consider the other problems through the prism of condensed matter physics, including the black holes and Planck constant.
@@ I Introduction
In Section V we consider the dynamics of de Sitter state which may follow from its thermodynamics. In the suggested phenomenological approach, it is assumed that the equations are modified in such a way, that away from equilibrium the gravitational dark matter behaves as the real stiff matter. The latter experiences the cool-down due to expansion, which is partially compensated by the energy exchange with the dark energy component. The energy exchange leads to the loss of the dark energy, and as a result both dark components have the power-law decay. Both components reach at present time the correct order of magnitude. How this phenomenology is supported by microscopic theory is an open question.
@@
In Section IX we discuss the dynamic origin of Planck constants $\hbar$ and $h=\hbar c$ as elements of the tetrads in Minkowski spacetime. We also discuss the space-time dimensions of the physical quantities. In particular the Planck constants $\hbar$ and /h/\!\!\!h have dimensions of time and space correspondingly, and thus they cannot be considered as the fundamental constants. The effective Planck constant $h_{\rm ac}$ emerging in the acoustic gravity in superfluid 4He has dimension of length and the size of the interatomic distance. Some dimensionless quantities can be quantized, which is considered on example of the entropies of the cosmological and black hole horizons.
@@ II.1 River model of quantum vacuum
In GR, the flow metric with acoustic horizon corresponds to the Painlevé-Gullstrand (PG) metric. Painleve ; Gullstrand This suggested that the gravitational field of the black hole with mass $M$ can be considered as the result of the flow of the quantum vacuum with superfluid velocity $v_{\rm s}^{2}=2GM/r$ with the event horizon at $v_{s}^{2}=c^{2}$. It is important that the PG metric is continuous across the horizon, which allows to study the thermodynamics of the vacua with horizons. Both for the real black hole and for its condensed matter analog, the Hawking radiation can be considered as semiclassical quantum tunneling across the horizon, see Ref. Volovik1999 for the analog black hole and Refs. ParikhWilczek2000 ; Srinivasan1999 for the real black hole.
@@ V.1 Phenomenological theory of two-component dynamics
The two-fluid thermodynamics of the de Sitter state discussed in Sec. IV is valid in thermal equilibrium, when $T=H/\pi$. However, the equilibrium state is impossible since the dark matter represents the heat bath for matter, which leads to the thermal nucleation of matter and thus to decay of the vacuum energy. Here we will not go into the details of this decay process, which depends on many different factors. We shall use the phenomenological approach suggested in Ref.Volovik2024 and apply to the Universe with the cosmological stiff matter. We assume that the Einstein equations are modified in such a way that the gravitational dark matter behaves as the real Zel’dovich stiff matter also in dynamics. Then the dark matter decays with time due to expansion, but it also gets energy from the energy exchange with the dark energy. On the other hand this energy exchange leads to decrease of the vacuum energy density, and the Universe finally approaches the equilibrium Minkowski vacuum state.
@@ VI.2 Hawking radiation as quantum tunneling
Moreover, the macroscopic quantum tunneling from the black hole to the white hole with the same mass demonstrates the curious thermodynamic properties of the white hole.Volovik2025P In particular, the entropy of the white hole is negative, $S_{\rm WH}(M)=-S_{\rm BH}(M)$. This is a consequence of a special type of event horizon that also occurs in a collapsing de Sitter universe with $H<0$. It is unclear whether negative entropy can arise in systems without an event horizon and in condensed matter analogues.
@@ VI.3 Emission of black holes as macroscopic quantum tunneling
The emitted black hole can be thought of as a type of particle, and the emission process contains a similar element of back reaction.Volovik2022G However, the ”particle” emitted by the black hole has a non-zero entropy. As a result, unlike the emission of a point particle, the emission rate of a small black hole increases by the entropy of the emitted black hole.HawkingHorowitz1995
@@ VI.4 Non-extensive entropy of black hole
The entropy of the conventional thermodynamic systems is the extensive. The entropy there is proportional to the volume of the system, and thus the splitting of the system with volume $V$ in two parts with volumes $V_{1}+V_{2}=V$ does not change the total entropy of the system, $S(V_{1}+V_{2})=S(V_{1})+S(V_{2})$, or $S(A,B)=S(A)+S(B)$. The black hole entropy does not satisfy the additivity condition. Instead one has the following non-additive composition rule for the black hole entropies:
@@ VII.1 Schwinger vs Unruh
If one tries to make the direct analogy between these processes, this is already problematic. The original state is the vacuum in the constant electric field. Being the vacuum it does not provide any physical acceleration. Acceleration in electric field appears only in the presence of a charged particle. That is why one can try to find the situation, when the two effects are physically connected. The connection may arise if we split the pair creation in several steps. In the first step the pair of particles is created by Schwinger mechanism and then in the further steps the Unruh process enters, which is caused by the acceleration of the created charged particles by electric field. These quantum processes should take place in unison, i.e. coherently. This means that in the quantum tunneling picture Volovik1999 ; ParikhWilczek2000 the corresponding exponents are multiplied.
@@ VIII.1 Vacuum energy in condensed matter
In condensed matter systems, such as superfluid 4He and superfluid phases of 3He, we know physics both at the ultraviolet (atomic) scale and in the infrared (hydrodynamic) limit. In this low-energy limit, physics is described by collective variables and quasiparticles, which form the analogue of the Universe. The effective quantum fields in these effective Universes have zero-point energies. These energies are formally diverging (in the same way as in our Universe) and require the ultraviolet cut-off. However, looking from the microscopic (atomic ultraviolet) side we can see that all these divergencies are formal and have no physical consequences. They do not determine the energy of the ground state of the quantum liquid – the energy of the ”quantum vacuum”, which is zero in equilibrium.
@@ VIII.2 Cosmological constant is zero in equilibrium
Eq.(69) is the general property of any vacuum state in any quantum system including the quantum vacuum of our Universe. It is important that it is $\epsilon_{\rm vac}$, and not $\epsilon$, that enters into Einstein’s equation as the cosmological constant $\Lambda$. This is shown using the so-called $q$-theory,KlinkhamerVolovik2008a ; KlinkhamerVolovik2008b ; KlinkhamerVolovik2008c where the physical vacuum is described by the 4-form field – the dynamical variable introduced by Hawking.Hawking1984+Duff+Wu From the equations for $q$-field and Einstein equation one obtains:
@@ VIII.3 Vacuum decay without fine-tuning
However, if the initial conditions are different, i.e. $\mu\neq m_{0}$, then from the equations of dynamics (again without dissipation) it follows that the Universe relaxes with oscillations to de Sitter attractor instead of the Minkowski vacuum state. This demonstrates that the special choice of initial conditions, $\mu=\mu_{0}$, is a kind of fine-tuning. The fine-tuning is also present in the Starobinsky inflation,StarobinskyYokoyama1994 ; Kofman1997 due to the special choice of the parameters of system, which leads to the Minkowski attractor. The Pauli-Zeldovich mechanism discussed by Kamenshchik and Starobinsky KamenshchikStarobinsky2018 also relies on the fine-tuning – the exact cancellation of contributions of relativistic bosons and relativistic fermions to the vacuum energy. In general case, i.e. without fine-tuning, the de Sitter attractor is inevitable. This reflects a special symmetry of the de Sitter state that puts this state on the same level as the Minkowski vacuum.
@@ IX.1 Composite tetrads from relative symmetry breaking
There are various scenarios for the emergence of gravity from more fundamental fields, such as quantum fermion fields. In particular, gravity can be constructed using the bilinear combinations of the fermionic fields: Akama1978 ; Diakonov2011 ; VladimirovDiakonov2012 ; VladimirovDiakonov2014 ; ObukhovHehl2012 ; Maiezza2022 ; Maitiniyazi2025
@@ IX.2 Dimensionful metric and dimensionless interval
and thus in this scenario of quantum gravity the metric is the fermionic quartet. In principle the signature can be also the dynamical variable $O_{ab}$,BondarenkoZubkov2022 ; Bondarenko2022 and $\eta_{ab}$ may also emerge as the vacuum expectation value of the corresponding symmetry breaking phase transition, $\eta_{ab}=<O_{ab}>$.
@@ IX.3 Dimensions of gauge field and mass
Such action can be extended to the objects of higher dimensions, which interact with the corresponding gauge fields: $1+1$ strings interacting with 2-form gauge field, $2+1$ branes interacting with the 3-form field and also $3+1$ medium interacting with the 4-form field.
@@
Since both the interval $ds$ and the action $S$ are dimensionless, from equation (77) it follows that the particle mass $M$ is also dimensionless, $M]=[S]=[s]=[L]^{0}=[t]^{0}=[1]$. Note that we discuss here the space and time dimensions of the physical quantities. In quantum field theories, the mass dimensions is also used, which will be discussed in Sec. IX.5.
@@ IX.4 Two Planck constants
Comparing the gradient term and the mass terms, and using the dimension of the metric, one again obtains that the mass $M$ is dimensionless, $[M]=[1]$. Then since the action $S$ and volume element $d^{4}x\,\sqrt{-g}$ are dimensionless, it follows that the scalar field is also dimensionless, $\Phi]^{2}=[M]=[S]=[1]$.
@@
This shows that the Minkowski elements of metric play the role of the Planck constants $\hbar$ and $h\equiv\hbar c$, which connect the energy and momentum of particles with the frequency and the wave vector. This suggests, that we have the following relations between the Planck constants and the Minkowski metric:
$g^{00}_{\rm Mink}\equiv\hbar^{2}\,\,,\,\,g^{ik}_{\rm Mink}\equiv/\!\!\!h^{2}\delta^{ik}\,.$ (83)
Both Planck constants, bar $\hbar$ and slash /h/\!\!\!h, are the elements of the metric and tetrads in Minkowski vacuum. The bar $\hbar$ is the time component of the inverse tetrad and has dimension of time, $\hbar]=[t]$ while the slash /h/\!\!\!h enters the space components of tetrads and has dimension of length, [/h]=[L][/\!\!\!h]=[L]. The presence of the Planck constants in the metric elements demonstrates that the metric field describes the dynamics of the quantum vacuum, rather than the space-time geometry. This is natural, since space-time itself with its metric arises from the dynamics of quantum fields. The speed of light, as an element of geometry, arises as the ratio of two Planck constants:
$c^{2}=\frac{/\!\!\!h^{2}}{\hbar^{2}}\,.$ (84)
In principle, there can be more than two Planck constants. For example, two Planck constants $\hbar$ and $\hbar_{g}$ were introduced in Ref. Terno2024 for matter and gravity correspondingly (see also Caro1999 ). This suggests that gravity will be more classical than matter if $\hbar_{g}\ll\hbar$. With slash component $h_{g}$ added one obtains four different Planck constants: $\hbar$, /h/\!\!\!h, $\hbar_{g}$ and slash $h_{g}$. Then in principle one may have two speeds of light: $c=/\!\!\!h/\hbar$ and $c_{g}=/\!\!\!h_{g}/\hbar_{g}$, although $c_{g}=c$ is more natural.
@@ IX.5 Planck constants do not enter diffeomorphism invariant expressions
Since $M$ is spacetime dimensionless, from equation $M=\hbar\omega$ it follows that the Planck constant has dimension of time, $\hbar]=[t]$, and the slash Planck constant has dimension of length, [/h]=[L][/\!\!\!h]=[L]. Since these ”constants” are not diffeomorphism invariant, they cannot enter the diffeomorphism invariant expressions. Example is the original action (72), which does not contain the Planck constants $\hbar$ and /h/\!\!\!h. This is the property of any action, if it is written in the diffeomorphism-invariant form. Let us consider several examples, when the Planck constants are removed from the expressions rewritten in diffeomorphism-invariant form.
@@
Since the potential $U$ has dimension of mass $M$, and thus is the space-time dimensionless, one obtains that $\bar{G}$} has the same dimension of length as /h/\!\!\!h, i.e. ${\bar{G}}]=[{/\!\!\!h}]=[L]$. This demonstrates that both $\bar{G}$} and /h/\!\!\!h are not diffeomorphism invariant, and thus they cannot be the fundamental constants. In the gravitational action they compensate each other:
$S=K\int d^{4}x\,\sqrt{-g}\,{\cal R}\,\,,\,\,K=\frac{1}{16\pi}\frac{/\!\!\!h}{{\bar{G}}}\equiv\frac{M_{\rm P}^{2}}{2}\,.$ (86)
@@
In Eq.(88), the fully dimensionless parameter $\alpha$ (the fine structure constant $\alpha=e^{2}//\!\!\!h$) connects the space-time dimensionless quantities: the electron rest energy $M_{e}$ and its correction $\Delta M_{e}$.
@@
In principle, some dimensionless parameters can be quantized. The symmetry numbers and topological charges are certainly quantized. The other possible example is the entropy of the black hole or the entropy of the Hubble volume, which are related to the event horizons.
@@
$E_{H}=\epsilon_{\rm vac}V_{H}=\frac{1}{2\hbar H}\frac{/\!\!\!h}{\bar{G}}=\frac{4M_{\rm P}}{T}\,M_{\rm P}\equiv NM_{\rm P}\,.$ (89)
@@
$S_{H}=\frac{A}{4/\!\!\!h\bar{G}}=\frac{\pi}{(\hbar H)^{2}}\frac{/\!\!\!h}{\bar{G}}=8\frac{M_{\rm P}^{2}}{T^{2}}=\frac{1}{2}N^{2}\,.$ (90)

### v11

@@ Abstract
Condensed matter analogs are useful when considering phenomena associated with the quantum vacuum. This is because in condensed matter, we know physics both in the infrared and in the ultraviolet limits, whereas in particle physics and gravity, physics on the trans-Planckian scale is unknown. One of the cornerstones of the connection between non-relativistic condensed matter and modern relativistic theories is the two-fluid hydrodynamics of superfluid helium, developed by Landau and Khalatnikov (Isaak Markovich Khalatnikov later founded and headed the Landau Institute). The dynamics and thermodynamics of the de Sitter state of the expansion of the Universe bear some features of a multi-fluid system. There are actually three components: the quantum vacuum, the gravitational component and relativistic matter. The expanding de Sitter vacuum serves as a thermal bath with local temperature, which is twice the Gibbons-Hawking temperature associated with the temperature of the cosmological horizon. This local temperature leads to the heating of the matter component and the gravitational component, which behaves as Zel’dovich stiff matter and represents dark matter. In equilibrium and in the absence of ordinary matter, the positive partial pressure of the dark matter compensates the negative partial pressure of the quantum vacuum. That is why, in full equilibrium, the total pressure is zero. This is rather similar to the correspondingly superfluid and normal components of superfluid liquid, which together produce a zero pressure of the liquid in the absence of environment. We propose the phenomenological theory, which describes the dynamics of dark energy and dark matter. If one assumes that, in the dynamics, gravitational dark matter behaves like real Zel’dovich stiff matter, it is shown that both components undergo power law decay due to the energy exchange between these components. It then follows that their values at the present time are of the correct order of magnitude. We also consider other problems through the prism of condensed matter physics, including black holes and the Planck constant.
@@ I Introduction
In Section V we consider the dynamics of de Sitter state which may follow from its thermodynamics. In the suggested phenomenological approach, it is assumed that the equations are modified in such a way, that away from equilibrium, the gravitational dark matter behaves as the real stiff matter. The latter experiences the cool-down due to expansion, which is partially compensated by the energy exchange with the dark energy component. The energy exchange leads to the loss of the dark energy, and as a result both dark components have the power-law decay. Both components reach at present time the correct order of magnitude. How this phenomenology is supported by microscopic theory is an open question.
@@
In Section IX we discuss the dynamic origin of Planck constants $\hbar$ and $\penalty 50\!\!\!h=\hbar c$ as elements of the tetrads in Minkowski spacetime. We also discuss the space-time dimensions of the physical quantities. In particular the Planck constants $\hbar$ and $\penalty 50\!\!\!h$ have dimensions of time and space correspondingly, and thus they cannot be considered as the fundamental constants. The effective Planck constant $\penalty 50\!\!\!h_{\rm ac}$ emerging in the acoustic gravity in superfluid 4He has dimension of length and the size of the interatomic distance. Some dimensionless quantities can be quantized, which is considered on example of the entropies of the cosmological and black hole horizons.
@@ II.1 River model of quantum vacuum
In general relativity (GR), the flow metric with acoustic horizon corresponds to the Painlevé-Gullstrand (PG) metric. Painleve ; Gullstrand This suggested that the gravitational field of the black hole with mass $M$ can be considered as the result of the flow of the quantum vacuum with superfluid velocity $v_{\rm s}^{2}=2GM/r$ with the event horizon at $v_{s}^{2}=c^{2}$. It is important that the PG metric is continuous across the horizon, which allows to study the thermodynamics of the vacua with horizons. Both for the real black hole and for its condensed matter analog, the Hawking radiation can be considered as semiclassical quantum tunneling across the horizon, see Ref. Volovik1999 for the analog black hole and Refs. ParikhWilczek2000 ; Srinivasan1999 for the real black hole.
@@ V.1 Phenomenological theory of two-component dynamics
The two-fluid thermodynamics of the de Sitter state discussed in Sec. IV is valid in thermal equilibrium, when $T=H/\pi$. However, the equilibrium state is impossible since the gravitational dark matter represents the heat bath for the ordinary matter, which leads to the thermal nucleation of matter and thus to decay of the vacuum energy. Here we will not go into the details of this decay process, which depends on many different factors. We shall use the phenomenological approach suggested in Ref.Volovik2024 and apply to the Universe with the cosmological stiff matter. We assume that the Einstein equations are modified in such a way that the gravitational dark matter behaves as the real Zel’dovich stiff matter also in dynamics. Then the dark matter decays with time due to expansion, but it also gets energy from the energy exchange with the dark energy. On the other hand this energy exchange leads to decrease of the vacuum energy density, and the Universe finally approaches the equilibrium Minkowski vacuum state.
@@ VI.2 Hawking radiation as quantum tunneling
Moreover, the macroscopic quantum tunneling from the black hole to the white hole with the same mass demonstrates the curious thermodynamic properties of the white hole.Volovik2025P In particular, the entropy of the white hole is negative, $S_{\rm WH}(M)=-S_{\rm BH}(M)$. This is a consequence of a special type of the white hole event horizon that also occurs in a collapsing de Sitter universe with $H<0$. It is unclear whether negative entropy can arise in systems without an event horizon and in condensed matter analogues.
@@ VI.3 Emission of black holes as macroscopic quantum tunneling
The emitted black hole can be thought of as a type of particle, and the emission process contains a similar element of back reaction.Volovik2022G However, such ”particle” emitted by the black hole has a non-zero entropy. As a result, unlike the emission of a point particle, the emission rate of a small black hole increases by the entropy of the emitted black hole.HawkingHorowitz1995
@@ VI.4 Non-extensive entropy of black hole
The entropy of the conventional thermodynamic systems is extensive. The entropy there is proportional to the volume of the system, and thus the splitting of the system with volume $V$ in two parts with volumes $V_{1}+V_{2}=V$ does not change the total entropy of the system, $S(V_{1}+V_{2})=S(V_{1})+S(V_{2})$, or $S(A,B)=S(A)+S(B)$.
The black hole entropy does not satisfy the additivity condition. Instead one has the following non-additive composition rule for the black hole entropies:
@@ VII.1 Schwinger vs Unruh
If one tries to make the direct analogy between the Schwinger and Unruh processes, this is already problematic. The original state is the vacuum in the constant electric field. Being the vacuum it does not provide any physical acceleration. Acceleration in electric field appears only in the presence of a charged particle. That is why one can try to find the situation, when the two effects are physically connected. The connection may arise if we split the pair creation in several steps. In the first step the pair of particles is created by Schwinger mechanism and then in the further steps the Unruh process enters, which is caused by the acceleration of the created charged particles by electric field. These quantum processes should take place in unison, i.e. coherently. This means that in the quantum tunneling picture,Volovik1999 ; ParikhWilczek2000 the corresponding exponents are added.
@@ VIII.1 Vacuum energy in condensed matter
In condensed matter systems, such as superfluid 4He and superfluid phases of 3He, we know physics both at the ultraviolet (atomic) scale and in the infrared (hydrodynamic) limit. In this low-energy limit, physics is described by collective variables and quasiparticles forming the analogue of the Universe. The effective quantum fields in these effective Universes have zero-point energies. These energies are formally diverging (in the same way as in our Universe) and require the ultraviolet cut-off. However, looking from the microscopic (atomic ultraviolet) side we can see that all these divergencies are formal and have no physical consequences. They do not determine the energy of the ground state of the quantum liquid – the energy of the ”quantum vacuum”, which is zero in equilibrium.
@@ VIII.2 Cosmological constant is zero in equilibrium
Eq.(69) is the general property of any vacuum state in any quantum system including the quantum vacuum of our Universe. It is important that it is $\epsilon_{\rm vac}$, and not $\epsilon$, that enters into Einstein’s equation as the cosmological constant $\Lambda$. This is shown using the so-called $q$-theory,KlinkhamerVolovik2008a ; KlinkhamerVolovik2008b ; KlinkhamerVolovik2008c where the physical vacuum is described by the 4-form field – the dynamical variable introduced by Hawking.Hawking1984 ; Duff+Wu From the equations for $q$-field and Einstein equation one obtains:
@@ VIII.3 Vacuum decay without fine-tuning
However, if the initial conditions are different, i.e. $\mu\neq\mu_{0}$, then from the equations of dynamics (again without dissipation) it follows that the Universe relaxes with oscillations to de Sitter attractor instead of the Minkowski vacuum state. This demonstrates that the special choice of initial conditions, $\mu=\mu_{0}$, is a kind of fine-tuning. The fine-tuning is also present in the Starobinsky inflation,StarobinskyYokoyama1994 ; Kofman1997 due to the special choice of the parameters of system, which leads to the Minkowski attractor. The Pauli-Zeldovich mechanism discussed by Kamenshchik and Starobinsky KamenshchikStarobinsky2018 also relies on the fine-tuning – the exact cancellation of contributions of relativistic bosons and relativistic fermions to the vacuum energy. In general case, i.e. without fine-tuning, the de Sitter attractor is inevitable. This reflects a special symmetry of the de Sitter state that puts this state on the same level as the Minkowski vacuum.
@@ IX.1 Composite tetrads from relative symmetry breaking
There are various scenarios for the emergence of gravity from more fundamental fields, such as quantum fermionic fields. In particular, gravity can be constructed using the bilinear combinations of the fermionic fields: Akama1978 ; Diakonov2011 ; VladimirovDiakonov2012 ; VladimirovDiakonov2014 ; ObukhovHehl2012 ; Maiezza2022 ; Maitiniyazi2025
@@ IX.2 Dimensionful metric and dimensionless interval
and thus in this scenario of quantum gravity the metric is the fermionic quartet. In principle, the signature of $\eta_{ab}$ can also come from the dynamical variable $O_{ab}$,BondarenkoZubkov2022 ; Bondarenko2022 if $\eta_{ab}$ emerges as the vacuum expectation value of the corresponding symmetry breaking phase transition, $\eta_{ab}=<O_{ab}>$.
@@ IX.3 Dimensions of gauge field and mass
Such action can be extended to the objects of higher dimensions, which interact with the corresponding gauge fields: $1+1$ strings interacting with 2-form gauge field, $2+1$ branes interacting with the 3-form field and also $3+1$ objects interacting with the 4-form field.
@@
Since both the interval $ds$ and the action $S$ are dimensionless, from equation (77) it follows that the particle mass $M$ is also dimensionless, $M]=[S]=[s]=[L]^{0}=[t]^{0}=[1]$. Note that we discuss here the space and time dimensions of the physical quantities. In quantum field theories, the mass dimension is also used, which will be discussed in Sec. IX.5.
@@ IX.4 Two Planck constants
Comparing the gradient term and the mass term, and using the dimension of the metric, one again obtains that the mass $M$ is dimensionless, $[M]=[1]$. Then since the action $S$ and volume element $d^{4}x\,\sqrt{-g}$ are dimensionless, it follows that the scalar field is also dimensionless, $\Phi]^{2}=[M]=[S]=[1]$.
@@
This shows that the Minkowski elements of metric play the role of the Planck constants $\hbar$ and $\penalty 50\!\!\!h\equiv\hbar c$, which connect the energy and momentum of particles with the frequency and the wave vector. This suggests, that we have the following relations between the Planck constants and the Minkowski metric:
$g^{00}_{\rm Mink}\equiv\hbar^{2}\,\,,\,\,g^{ik}_{\rm Mink}\equiv/\penalty 50\!\!\!h^{2}\delta^{ik}\,.$ (83)
Both Planck constants, bar $\hbar$ and slash $\penalty 50\!\!\!h$, are the elements of the metric and tetrads in Minkowski vacuum. The bar $\hbar$ is the time component of the inverse tetrad and has dimension of time, $\hbar]=[t]$, while the slash $\penalty 50\!\!\!h$ enters the space components of tetrads and has dimension of length, $\penalty 50\!\!\!h]=[L]$. The presence of the Planck constants in the metric elements demonstrates that the metric field describes the dynamics of the quantum vacuum, rather than the space-time geometry. This is natural, since space-time itself with its metric arises from the dynamics of quantum fields. The speed of light, as an element of geometry, arises as the ratio of two Planck constants:
$c^{2}=\frac{/\penalty 50\!\!\!h^{2}}{\hbar^{2}}\,.$ (84)
In principle, there can be more than two Planck constants. For example, two Planck constants $\hbar$ and $\hbar_{g}$ were introduced in Ref. Terno2024 for matter and gravity correspondingly (see also Caro1999 ). This suggests that gravity will be more classical than matter if $\hbar_{g}\ll\hbar$. With slash component $\penalty 50\!\!\!h_{g}$ added one obtains four different Planck constants: $\hbar$, $\penalty 50\!\!\!h$, $\hbar_{g}$ and slash $\penalty 50\!\!\!h_{g}$. Then in principle one may have two speeds of light: $c=/\penalty 50\!\!\!h/\hbar$ and $c_{g}=/\penalty 50\!\!\!h_{g}/\hbar_{g}$, although $c_{g}=c$ is more natural.
@@ IX.5 Planck constants do not enter diffeomorphism invariant expressions
Since $M$ is spacetime dimensionless, from equation $M=\hbar\omega$ it follows that the Planck constant has dimension of time, $\hbar]=[t]$, and the slash Planck constant has dimension of length, $\penalty 50\!\!\!h]=[L]$. Since these ”constants” are not diffeomorphism invariant, they cannot enter the diffeomorphism invariant expressions. Example is the original action (72), which does not contain the Planck constants $\hbar$ and $\penalty 50\!\!\!h$. This is the property of any action, if it is written in the diffeomorphism-invariant form. Let us consider several examples, when the Planck constants are removed from the expressions rewritten in diffeomorphism-invariant form.
@@
Since the potential $U$ has dimension of mass $M$, and thus is the space-time dimensionless, one obtains that $\bar{G}$} has the same dimension of length as $\penalty 50\!\!\!h$, i.e. ${\bar{G}}]=[{/\penalty 50\!\!\!h}]=[L]$. This demonstrates that both $\bar{G}$} and $\penalty 50\!\!\!h$ are not diffeomorphism invariant, and thus they cannot be the fundamental constants. In the gravitational action they compensate each other:
$S=K\int d^{4}x\,\sqrt{-g}\,{\cal R}\,\,,\,\,K=\frac{1}{16\pi}\frac{/\penalty 50\!\!\!h}{{\bar{G}}}\equiv\frac{M_{\rm P}^{2}}{2}\,.$ (86)
@@
In Eq.(88), the fully dimensionless parameter $\alpha$ (the fine structure constant $\alpha=e^{2}//\penalty 50\!\!\!h$) connects the space-time dimensionless quantities: the electron rest energy $M_{e}$ and its correction $\Delta M_{e}$.
@@
In principle, some dimensionless parameters can be quantized. The numbers related to symmetry and topological charges are certainly quantized. The other possible example is the entropy of the black hole or the entropy of the Hubble volume, which are related to the event horizons.
@@
$E_{H}=\epsilon_{\rm vac}V_{H}=\frac{1}{2\hbar H}\frac{/\penalty 50\!\!\!h}{\bar{G}}=\frac{4M_{\rm P}}{T}\,M_{\rm P}\equiv NM_{\rm P}\,.$ (89)
@@
$S_{H}=\frac{A}{4/\penalty 50\!\!\!h\bar{G}}=\frac{\pi}{(\hbar H)^{2}}\frac{/\penalty 50\!\!\!h}{\bar{G}}=8\frac{M_{\rm P}^{2}}{T^{2}}=\frac{1}{2}N^{2}\,.$ (90)